# Higher property T and below-rank phenomena of lattices

### v1

@@ 2.11. Group algebra reformulation of higher T
In this subsection we discuss group algebra criteria for higher property T. We fix a free resolution $P_{\mathbin{{\bullet}}}$ as in §2.2. We may and will assume that $P_{0}\cong{\mathbb{C}}\Gamma$. Then the Laplace operator $\Delta_{0}\colon P_{0}\to P_{0}$ is given by an element in the group algebra ${\mathbb{C}}\Gamma$. The motivation for our discussion is the celebrated Ozawa criterion.
Theorem 48 ([OzawaT]).
@@
A criterion of a similar spirit was given in [Bader-Nowak] for $T_{n}]$, with a caveat - the implication (1) $\Rightarrow$ (7) of Lemma 35 was yet not observed at the time of its publication. In retrospect, we do get the following.
Theorem 49 ([Bader-Nowak]).
@@
By Lemma 35, the group $\Gamma$ has $T_{n}]$ if and only if $\Delta_{k}$ is invertible as a matrix over $C^{*}\Gamma$ for every $1\leq k\leq n$. Hence the theorem follows from [Bader-Nowak, Proposition 16]. ∎
@@
In the upcoming discussion, towards the proof of Theorem 10, we will freely use the terminology introduced in [Bader-Nowak].
@@
If we start with a free ${\mathbb{Q}}\Gamma$-resolution $P_{\mathbin{{\bullet}}}$, then the Laplace operators $\Delta_{k}$ are represented by matrices over ${\mathbb{Q}}\Gamma$. In Theorems 48 and 10, the matrix entries of the elements $x_{i}$ can then be chosen to be in ${\mathbb{Q}}\Gamma$ (cf. [Bader-Nowak]*Corollary 15).
@@
Let $A$ be a unital *-algebra that is archimedean has a trivial infinitesimal ideal. In particular, $A$ injects into its C*-envelope $C=C^{*}(A)$ [Bader-Nowak]*§2.2. Then every intermediate *-algebra $A\subset B\subset C$ is also archimedean and has a trivial infinitesimal ideal. We may identify naturally $C^{*}(B)=C$.
@@
The entire lemma follows from its last sentence. [Bader-Nowak, Proposition 9] gives the last sentence of the Lemma for $B=C$. In general, $A_{+}\subseteq B_{+}\subseteq C_{+}$ and the last sentence follows from the case $B=C$. ∎
@@
First note that $\Delta_{0}(\Delta_{0}-\epsilon)\Delta_{0}\in{\mathbb{C}}\Gamma_{+}$ implies that every unitary $\Gamma$-representation $V$ with $V^{\Gamma}=0$ has a spectral gap and $\Gamma$ has property T. So we may and will assume that $\Gamma$ has property T. We denote by $z$ the Kazhdan projection in $C^{*}\Gamma$ and set e=1ze=1-z. Then $\Delta_{0}e=\Delta_{0}$ and it is invertible in $C_{0}^{*}\Gamma=eC^{*}\Gamma$. By Lemma 35 $\Gamma$ has $(T_{n})$ if and only if $\Delta_{k}$ is invertible as a matrix over $C_{0}^{*}\Gamma$ for every $1\leq k\leq n$. Thus the theorem follows from Lemma 52 with regard to the following setting: We take $A=M_{m_{k}}({\mathbb{C}}\Gamma)$, identify $C^{*}(A)=M_{m_{k}}(C^{*}\Gamma)$, view $C^{*}\Gamma$ as a unital subalgebra of $C^{*}(A)$ and $e$ as a central idempotent of $C^{*}(A)$. Moreover, we take $a_{0}=\Delta_{0}$ and $a=\Delta_{k}$. Compare [Bader-Nowak, §2.5]. ∎
@@ 2.12.1. Products
and its associated long exact sequence in bounded cohomology [frigerio]*Section 1.5. This long exact sequence starts as
@@ 2.12.2. Quotients
While property T passes to quotients, this is not the case with the higher analogues. For example, uniform lattices in the exceptional simple Lie group $F_{4}^{(-20)}$ have $T_{2}]$, but Francesco Fournier-Facio has verified that they admit quotients that do not satisfy $(T_{2})$.
@@ 2.12.3. Subgroups of finite index
The other direction is less obvious. The following result is [badsau]*Theorem 6.21.
@@
By same result from [badsau] property $(T_{n})$ for a group $G$ implies that the cohomology of every subgroup of $G$ of finite index with coefficients in a weakly mixing unitary representation vanishes in degrees up to $n$. But is less clear that property $(T_{n})$ passes to subgroups of finite index in general.
@@ 2.13. Back to the finiteness assumption
It follows that for a property T group $\Gamma$, and for the corresponding unital C*-algebra $C_{0}^{*}\Gamma$, without any finiteness assumptions, the property of having for every $k\leq n$, $H_{k}(\Gamma,C_{0}^{*}\Gamma)=0$, implies $(T_{n})$ and the entire variety of properties listed in Theorem 11 and Theorem 12. This property is equivalent to $(T_{n})$ under $FP_{\infty}({\mathbb{Q}})$, but this equivalence is not known in general, which brings us back to Question 15.
@@ 3. Lattices in semisimple groups
Finally, in §3.4 we isolate low‑degree consequences in degrees 1 and 2, organizing a web of conjectures around two spectral gap conjectures (Conjecture 97 and 98) and its implications for property $\tau$, coamenable subgroups, invariant random subgroup and character rigidity, and stability and approximation problems.
@@ 3.1. From Theorem 1 towards Conjecture 5 and Theorem 7
Theorem 1 was proved by the two authors of the present paper in [badsau]. The following is a rough sketch of the proof in [badsau].
@@
    Use results from representation theory by Zuckerman [Zuckerman] and Borel-Wallach [Borel-Wallach] (see also Vogan-Zuckerman [vogan+zuckerman]) saying that the cohomology of $G$ vanishes in degrees below the rank for non-trivial irreducible unitary representations.
@@
    Every unitary representation has a direct integral decomposition into irreducible unitary representations. The direct integral decomposition is compatible with reduced cohomology by [Blanc]*Théorème 7.2. As a consequence, one obtains the vanishing of the reduced cohomology of $G$ for all unitary representations without invariant vectors in degrees below the rank.
@@
    Finally, to deduce that a similar vanishing result applies to an arbitrary (maybe non-cocompact) lattice, a replacement for the Shapiro lemma is developed in [badsau] based on recent advances in geometric group theory. See the discussion after Theorem 77. This implies the same vanishing result for all lattices in $G$.
@@
We will start with the following remarkable theorem in degree $1$, which was conjectured in [BFGM]. It was proved recently by Oppenheim [Oppenheim] and de Laat–de la Salle [Sa-Le-at]; over a non-archimedean field it was proved earlier by Lafforgue [Lafforgue].
@@
In attempting a generalization to higher degrees, Conjecture 62 below was formulated in a recent joint work with Saar Bader and Shaked Bader [BBBS].
@@
The main work of [BBBS] lies in proving the following theorem, which relates Conjecture 62 and Conjecture 5.
Theorem 63 ([BBBS]).
@@
If one tries to apply the same proof strategy as for Theorem 1 sketched above, the first and second step immediately break down. There is no decomposition into irreducibles in the world of Banach space representations. A major achievement of [BBBS], in which we prove Theorem 63 and its consequences Theorem 7 and Corollary 8, is that we can avoid the results from Vogan-Zuckerman. This also yields a new proof of Theorem 1, different from [badsau].
Next we sketch the main ideas behind the proof of Theorem 63. For more details we direct the reader to the forthcoming paper [BBBS].
@@
The $n$-simplices of the opposition complex consist of ordered pairs of opposite $n$-simplices in the Tits building. The dimension of the opposition complex is thus the same as the one of the Tits building, namely r1r-1. There is an analogue of the Solomon-Tits theorem for the opposition complex saying that it is (r2)(r-2)-connected [heydebreck]. Further, the stabilizers of the $k$-dimensional simplices with respect to the natural $G$-action on the opposition complex are Levi subgroups with semisimple rank rk1r-k-1. The proof of the vanishing of cohomology part of Proposition 65 follows by mathematical induction using a cohomological induction argument. Finally, the Hausdorffness of the cohomology at the rank follows by an ultrapower argument. It is here, in taking an ultrapower, that Conjecture 62 is used to show that we gain no further invariants in the limiting process.
@@
For unitary representations, the conjecture follows by considering a subgroup of $G$ of the form $\operatorname{SL}_{2}(\mathbb{R})\ltimes E$, where $E$ is a finite dimensional representation of $\operatorname{SL}_{2}(\mathbb{R})$, and using its relative property T to find an $E$-fixed point, which gives a contradiction by the Howe-Moore theorem. For an equivalent norm on a Hilbert space, following an idea by Shalom, one may use unitarization with respect to amenable subgroups. This was carried out by Glasner and Gorfine and will appear in their forthcoming work [GG]. For the other spaces in the list above, the proof is by comparison to the unitary case, see [BFGM] and [Olivier]. We expect that the class above could be enlarged dramatically by the method of interpolation.
@@
However, the obvious analogue of Conjecture 62 does hold, due to [Olivier]. Translating the proof of Theorem 63 goes rather smoothly with only one crucial difference - when proving Proposition 66 using a version of Lemma 56, in showing that $H^{1}_{b}(C,B^{r}(S,V))=0$ we cannot use the fact that the space $B^{r}(S,V)$ is super-reflexive. Here $V$ denotes the coefficients in question. Rather, we prove that this space is L-embedded and we invoke the fixed point theorem of [BGM] to conclude the required cohomological vanishing. ∎
We end this section by remarking that the extra work in proving the $p=1$ case of Theorem 7 is quite rewarding - see §3.3.2 and §3.3.3 for geometric applications of its Corollary 8.
@@ 3.2.1. Gromov’s conjecture on $L^{p}$-cohomology
In this subsection we focus on specific Banach representations of a group $\Gamma$, namely the regular ones, $L^{p}(\Gamma)$, for $p\in[1,\infty)$. The resulting (continuous) group cohomology with coefficients in $L^{p}(\Gamma)$ is called the $L^{p}$-cohomology of $\Gamma$ – a subject that owes a lot to the work of Pansu and Gromov, see for instance [Pansu] and [Gromov93]*§8.
We remark that for $p=2$ the $L^{p}$-cohomology is a module over the von Neumann algebra $L\Gamma$ and the availability of Murray-von Neumann dimension is a valuable tool in this case. However, we are interested in arbitrary $p$ and it was shown in [Monodlp] that a reasonable dimension theory is not expected for $p>2$, so our discussion below is dimension free  [^11] Nonetheless, note that an entropy-type dimension for sofic groups and for $p<2$ was suggested in [Hayes1, Hayes2]..
The first $L^{p}$-cohomology is well studied. See [pansu-cohom, BP, Yu, Nica, BFGM, cornulier+tessera-cohom_banach, cornulier+tessera-contracting]. For every $p$, the first $L^{p}$-cohomology is Hausdorff if and only if $\Gamma$ is non-amenable. If $\Gamma$ has property T and $p\in[1,2]$, then $H^{1}(\Gamma,L^{p}(\Gamma))=0$ by [BFGM] (for $p>1$) and [BGM] (for $p=1$). A similar statement for property T groups in general fails for $p>2$. Pansu proved that a lattice in a Lie group of rank $1$ has $H^{1}(\Gamma,L^{p}(\Gamma))\neq 0$ for sufficiently large $p$ [pansu-cohom]. In particular, property $T_{2}]$, which is satisfied by $\operatorname{Sp}(n,1)$ and its cocompact lattices, is not robust enough to ensure that the $L^{p}$-cohomology vanishes for every $p\in[1,\infty)$. Later, Bourdon-Pajot generalized Pansu’s result to all non-elementary Gromov-hyperbolic groups [BP], see also [Yu, Nica]. The situation for a lattice in higher rank simple Lie groups is different: There we have $H^{1}(\Gamma,L^{p}(\Gamma))=0$ for every $p\in[1,\infty)$ by an unpublished result of Pansu and [BFGM, Theorem B].
The $L^{p}$-cohomology in higher degrees and for $p\neq 2$ is much less understood. Using Lemma 54, one easily sees that for a product of finite type non-amenable groups, $\Gamma=\Gamma_{1}\times\cdots\Gamma_{n}$, the $L^{p}$-cohomology vanishes in every degree $j<n$. It is also not hard to see that in degree $n$ it equals the $L^{p}$-tensor product of the first $L^{p}$-cohomologies of its factors. See [Gromov93, p. 252] for a de Rham theoretic proof. This fact motivated Gromov to conjecture a cohomological vanishing below-rank for symmetric spaces and buildings, see [Gromov93, p. 253]. Moreover, Gromov conjectured that at the rank the $L^{p}$-cohomology is Hausdorff and that it doesn’t vanish for $p$ large enough, in fact for all $p>1$ in the building case. This conjecture is now fully confirmed.
@@
Over archimedean fields, both the non-vanishing in Statement (2) and the vanishing in Statement (3) are due to Bourdon-Rémy [BR23]*Theorem A. Over non-archimedean fields, the non-vanishing in Statement (2) is due to [Lopez, Corollary 1.2] and the vanishing in Statement (3) follows by cohomological dimension. In both cases, and both for the groups and their lattices, Statement (1) and the Hausdorffness in Statement (2) is due to [BBBS], being a special case of Corollary 8. For cocompact lattices, the full result follows by Shapiro Lemma and $L^{p}$-induction. ∎
@@
For $\Gamma$ hyperbolic group, following the work of Bourdon-Pajot [BP], Nica constructed for every $p$ large enough a non-trivial element in $H^{1}(\Gamma,H^{1}(\Gamma,\ell^{p}\Gamma))$, see [Nica, Theorem 8]. Using the Lyndon-Hochschild-Serre spectral sequence we have the identification $H^{1}(\Gamma,H^{1}(\Gamma,\ell^{p}\Gamma))=H^{2}(\Gamma\times\Gamma,\ell^{p}\Gamma)$ and we interpret the Nica cocycle accordingly. In particular, we get that $H^{2}(\Gamma\times\Gamma,\ell^{p}\Gamma)\neq 0$.

### v2

@@ 2.11. Group algebra reformulation of higher T
In this subsection we discuss group algebra criteria for higher property T. We fix a free resolution $P_{\mathbin{{\bullet}}}$ as in §2.2. Recall that $P_{0}={\mathbb{C}}\Gamma$. Then the Laplace operator $\Delta_{0}\colon P_{0}\to P_{0}$ is given by an element in the group algebra ${\mathbb{C}}\Gamma$. The motivation for our discussion is the celebrated Ozawa criterion.
Theorem 54 ([83]).
@@
A criterion of a similar spirit was given in [13] for $T_{n}]$, with a caveat - the implication (1) $\Rightarrow$ (7) of Lemma 40 was yet not observed at the time of its publication. In retrospect, we do get the following.
Theorem 55 ([13]).
@@
By Lemma 40, the group $\Gamma$ has $T_{n}]$ if and only if $\Delta_{k}$ is invertible as a matrix over $C^{*}\Gamma$ for every $1\leq k\leq n$. Hence the theorem follows from [13, Proposition 16]. ∎
@@
In the upcoming discussion, towards the proof of Theorem 15, we will freely use the terminology introduced in [13].
@@
If we start with a free ${\mathbb{Q}}\Gamma$-resolution $P_{\mathbin{{\bullet}}}$, then the Laplace operators $\Delta_{k}$ are represented by matrices over ${\mathbb{Q}}\Gamma$. In Theorems 54 and 15, the matrix entries of the elements $x_{i}$ can then be chosen to be in ${\mathbb{Q}}\Gamma$ (cf. [13]*Corollary 15).
@@
Let $A$ be a unital *-algebra that is archimedean has a trivial infinitesimal ideal. In particular, $A$ injects into its C*-envelope $C=C^{*}(A)$ [13]*§2.2. Then every intermediate *-algebra $A\subset B\subset C$ is also archimedean and has a trivial infinitesimal ideal. We may identify naturally $C^{*}(B)=C$.
@@
The entire lemma follows from its last sentence. [13, Proposition 9] gives the last sentence of the Lemma for $B=C$. In general, $A_{+}\subseteq B_{+}\subseteq C_{+}$ and the last sentence follows from the case $B=C$. ∎
@@
First note that $\Delta_{0}(\Delta_{0}-\epsilon)\Delta_{0}\in{\mathbb{C}}\Gamma_{+}$ implies that every unitary $\Gamma$-representation $V$ with $V^{\Gamma}=0$ has a spectral gap and $\Gamma$ has property T. So we may and will assume that $\Gamma$ has property T. We denote by $z$ the Kazhdan projection in $C^{*}\Gamma$ and set e=1ze=1-z. Then $\Delta_{0}e=\Delta_{0}$ and it is invertible in $C_{0}^{*}\Gamma=eC^{*}\Gamma$. By Lemma 40 $\Gamma$ has $(T_{n})$ if and only if $\Delta_{k}$ is invertible as a matrix over $C_{0}^{*}\Gamma$ for every $1\leq k\leq n$. Thus the theorem follows from Lemma 58 with regard to the following setting: We take $A=M_{m_{k}}({\mathbb{C}}\Gamma)$, identify $C^{*}(A)=M_{m_{k}}(C^{*}\Gamma)$, view $C^{*}\Gamma$ as a unital subalgebra of $C^{*}(A)$ and $e$ as a central idempotent of $C^{*}(A)$. Moreover, we take $a_{0}=\Delta_{0}$ and $a=\Delta_{k}$. Compare [13, §2.5]. ∎
@@ 2.12.1. Products
and its associated long exact sequence in bounded cohomology [50]*Section 1.5. This long exact sequence starts as
@@ 2.12.2. Quotients
While property T passes to quotients, this is not the case with the higher analogues. For example, uniform lattices in the exceptional simple Lie group $F_{4}^{(-20)}$ have $(T_{3})$ (see Theorem 5) but Fournier-Facio has shown [46] that they admit quotients that do not satisfy $(T_{2})$.
@@ 2.12.3. Subgroups of finite index
The other direction is less obvious. The following result is [15]*Theorem 6.21.
@@
By same result from [15] property $(T_{n})$ for a group $G$ implies that the cohomology of every subgroup of $G$ of finite index with coefficients in a weakly mixing unitary representation vanishes in degrees up to $n$. But is less clear that property $(T_{n})$ passes to subgroups of finite index in general.
@@ 2.13. Back to the finiteness assumption
Assume now that $\Gamma$ is finitely generated, but do not assume any further finiteness properties. Note that Lemma 45 and Remark 46 still hold, thus we have that $H_{k}(\Gamma,C_{0}^{*}\Gamma)=0$ implies that $\Gamma$ has property T, thus the C*-algebra $C_{0}^{*}\Gamma$ is unital by Lemma 43 and Convention 44. It follows from Lemma 65 that under merely finite generation assumption, the property of having for every $k\leq n$, $H_{k}(\Gamma,C_{0}^{*}\Gamma)=0$, implies $(T_{n})$ and the entire variety of properties listed in Theorem 16 and Theorem 17. This property is equivalent to $(T_{n})$ under $FP_{\infty}({\mathbb{Q}})$, but this equivalence is not known in general, which brings us back to Question 20.
@@ 3. Lattices in semisimple groups
Finally, in §3.4 we isolate low‑degree consequences in degrees 1 and 2, organizing a web of conjectures around two spectral gap conjectures (Conjecture 105 and 106) and its implications for property $\tau$, coamenable subgroups, invariant random subgroup and character rigidity, and stability and approximation problems.
@@ 3.1. From Theorem 1 towards Conjecture 7 and Theorem 11
Theorem 1 was proved by the two authors of the present paper in [15]. The following is a rough sketch of the proof in [15].
@@
    Use results from representation theory by Zuckerman [101] and Borel-Wallach [22] (see also Vogan-Zuckerman [97]) saying that the cohomology of $G$ vanishes in degrees below the rank for non-trivial irreducible unitary representations.
@@
    Every unitary representation has a direct integral decomposition into irreducible unitary representations. The direct integral decomposition is compatible with reduced cohomology by [19]*Théorème 7.2. As a consequence, one obtains the vanishing of the reduced cohomology of $G$ for all unitary representations without invariant vectors in degrees below the rank.
@@
    Finally, to deduce that a similar vanishing result applies to an arbitrary (maybe non-cocompact) lattice, a replacement for the Shapiro lemma is developed in [15] based on recent advances in geometric group theory. See the discussion after Theorem 85. This implies the same vanishing result for all lattices in $G$.
@@
We will start with the following remarkable theorem in degree $1$, which was conjectured in [10]. It was proved recently by Oppenheim [81] and de Laat–de la Salle [37]; over a non-archimedean field it was proved earlier by Lafforgue [66].
@@
In attempting a generalization to higher degrees, Conjecture 68 below was formulated in a recent joint work with Saar Bader and Shaked Bader [7].
@@
The main work of [7] lies in proving the following theorem, which relates Conjecture 68 and Conjecture 7.
Theorem 69 ([7]).
@@
If one tries to apply the same proof strategy as for Theorem 1 sketched above, the first and second step break down. There is no decomposition into irreducibles in the world of Banach space representations. Note that the first and second step are easy and quite standard in the unitary case. The innovation of the proof of Theorem 1 lies mainly in steps (4), (5) and (6), which remain indispensable in the proof of Theorem 69. The major innovation of [7], in which we prove Theorem 69 and its consequences Theorem 11 and Corollary 12, is that we can avoid the results from Vogan-Zuckerman.
Next we sketch the main ideas behind the proof of Theorem 69. For more details we direct the reader to the forthcoming paper [7].
@@
The $n$-simplices of the opposition complex consist of ordered pairs of opposite $n$-simplices in the Tits building. The dimension of the opposition complex is thus the same as the one of the Tits building, namely r1r-1. There is an analogue of the Solomon-Tits theorem for the opposition complex saying that it is (r2)(r-2)-connected [98]. Further, the stabilizers of the $k$-dimensional simplices with respect to the natural $G$-action on the opposition complex are Levi subgroups with semisimple rank rk1r-k-1. The proof of the vanishing of cohomology part of Proposition 71 follows by mathematical induction using a cohomological induction argument. Finally, the Hausdorffness of the cohomology at the rank follows by an ultrapower argument. It is here, in taking an ultrapower, that Conjecture 68 is used to show that we gain no further invariants in the limiting process.
@@
For unitary representations, the conjecture follows by considering a subgroup of $G$ of the form $\operatorname{SL}_{2}(\mathbb{R})\ltimes E$, where $E$ is a finite dimensional representation of $\operatorname{SL}_{2}(\mathbb{R})$, and using its relative property T to find an $E$-fixed point, which gives a contradiction by the Howe-Moore theorem. For an equivalent norm on a Hilbert space, following an idea by Shalom, one may use unitarization with respect to amenable subgroups. This was carried out by Glasner and Gorfine and will appear in their forthcoming work [53]. For the other spaces in the list above, the proof is by comparison to the unitary case, see [10] and [80]. We expect that the class above could be enlarged dramatically by the method of interpolation.
@@
However, the obvious analogue of Conjecture 68 does hold, due to [80]. Translating the proof of Theorem 69 goes rather smoothly with only one crucial difference - when proving Proposition 72 using a version of Lemma 62, in showing that $H^{1}_{b}(C,B^{r}(S,V))=0$ we cannot use the fact that the space $B^{r}(S,V)$ is super-reflexive. Here $V$ denotes the coefficients in question. Rather, we prove that this space is L-embedded and we invoke the fixed point theorem of [8] to conclude the required cohomological vanishing. ∎
@@ 3.2. Cohomological below-rank phenomena
3.2.1. Borel’s stability and $K$-theory
Borel’s stability theorem asserts that, for a fixed degree, the rational cohomology of arithmetic groups such as $\operatorname{SL}_{n}({\mathbb{Z}})$ stabilizes as $n\to\infty$, and coincides with the corresponding stable cohomology of the associated real Lie group. Specifically, Borel proves the following result for $\operatorname{SL}_{n}$; similar results hold for other classical groups.
Theorem 75 (Borel [24]).
For every integer $i\geq 0$, the maps induced by the standard inclusions The restriction map
$H^{i}_{c}\bigl(\operatorname{SL}_{n}({\mathbb{R}}),{\mathbb{C}}\bigr)\to H^{i}\bigl(\operatorname{SL}_{n}({\mathbb{Z}}),{\mathbb{C}}\bigr)$
is an isomorphism for $0\leq i<(n-1)/4$.
Via this result, Borel identified the stable rational cohomology as an exterior algebra on explicit generators in odd degrees, now known as the Borel classes. This result plays a central role in the computation of rational algebraic $K$-theory of rings of integers: via the comparison between group cohomology and algebraic $K$-theory, Borel’s stability allows one to determine the ranks of $K_{i}({\mathbb{Z}})\otimes{\mathbb{Q}}$ and, more generally, of $K_{i}(\mathcal{O}_{F})\otimes{\mathbb{Q}}$ for number fields $F$.
Our Theorem 10 from [15] significantly improves Borel’s stability range and allows for arbitrary unitary coefficients.
This yields new computations of the low-degree cohomology of arithmetic groups; it does not yield new computations of algebraic $K$-theory as the slope of the stability range does not matter here. For the trivial representation and congruence lattices, Li-Sun independently provide a similar improvement of Borel’s stability range [69]*Theorem 1.8. Li-Sun’s paper relies heavily on the work of Franke [49], which plays no role in our approach.
The following result of Borel-Yang was the key to their solution of the rank conjecture in algebraic K-theory. The paper by Borel-Yang is heavily based on the work of Blasius-Franke-Grunewald [20], which itself relies on Franke’s work [49].
Theorem 76 (Borel-Yang [23]).
The restriction map
$H^{\ast}_{c}\bigl(\operatorname{SL}_{n}({\mathbb{R}}),{\mathbb{C}}\bigr)\to H^{\ast}\bigl(\operatorname{SL}_{n}({\mathbb{Q}}),{\mathbb{C}}\bigr)$
is an isomorphism in all degrees. More generally, if $k$ is a number field and $\mathbf{G}$ is a connected, simply connected, almost simple $k$-algebraic group, then the restriction map
$H_{c}^{*}\bigl(\mathbf{G}(k\otimes{\mathbb{R}}),{\mathbb{C}}\bigr)\to H^{*}\bigl(\mathbf{G}(k),{\mathbb{C}}\bigr)$
is an isomorphism in all degrees.
The correct generalization of Borel-Yang’s theorem to unitary coefficients is Theorem 89 in §3.3.1 and needs the adelic framework.
We refer to [15]*Theorems F and G for further generalizations of Borel-Yang’s theorem. In fact, the entire unitary cohomology theory of $\mathbf{G}(k)$ is determined by that of $\mathbf{G}(\mathbb{A}(k))$, at least when the $k$-rank of $\mathbf{G}$} is at least 2. Although ${\mathbf{G}}(k)$ is not of type I and a complete understanding of its unitary dual is out of reach, Theorem 7.9 and Corollary 7.10 of [15] determine the cohomological unitary dual of ${\mathbf{G}}(k)$ completely.
Also these results from [15] are proved using the strategy outlined in §3.1, with a heavy reliance on the recent advances in geometric group theory discussed in §3.3.1. We think it would be very interesting to understand better the connection of our methods to Franke’s work [49].
@@ 3.2.2. Gromov’s conjecture on $L^{p}$-cohomology
In this subsection we focus on specific Banach representations of a group $\Gamma$, namely the regular ones, $L^{p}(\Gamma)$, for $p\in[1,\infty)$. The resulting (continuous) group cohomology with coefficients in $L^{p}(\Gamma)$ is called the $L^{p}$-cohomology of $\Gamma$ – a subject that owes a lot to the work of Pansu and Gromov, see for instance [85] and [55]*§8.
We remark that for $p=2$ the $L^{p}$-cohomology is a module over the von Neumann algebra $L\Gamma$ and the availability of Murray-von Neumann dimension is a valuable tool in this case. However, we are interested in arbitrary $p$ and it was shown in [77] that a reasonable dimension theory is not expected for $p>2$, so our discussion below is dimension free  [^11] Nonetheless, note that an entropy-type dimension for sofic groups and for $p<2$ was suggested in [59, 60]..
The first $L^{p}$-cohomology is well studied. See [84, 25, 99, 78, 10, 34, 33]. For every $p$, the first $L^{p}$-cohomology is Hausdorff if and only if $\Gamma$ is non-amenable. If $\Gamma$ has property T and $p\in[1,2]$, then $H^{1}(\Gamma,L^{p}(\Gamma))=0$ by [10] (for $p>1$) and [8] (for $p=1$). A similar statement for property T groups in general fails for $p>2$. Pansu proved that a lattice in a Lie group of rank $1$ has $H^{1}(\Gamma,L^{p}(\Gamma))\neq 0$ for sufficiently large $p$ [84]. In particular, property $T_{2}]$, which is satisfied by $\operatorname{Sp}(n,1)$ and its cocompact lattices, is not robust enough to ensure that the $L^{p}$-cohomology vanishes for every $p\in[1,\infty)$. Later, Bourdon-Pajot generalized Pansu’s result to all non-elementary Gromov-hyperbolic groups [25], see also [99, 78]. The situation for a lattice in higher rank simple Lie groups is different: There we have $H^{1}(\Gamma,L^{p}(\Gamma))=0$ for every $p\in[1,\infty)$ by an unpublished result of Pansu and [10, Theorem B].
The $L^{p}$-cohomology in higher degrees and for $p\neq 2$ is much less understood. Using Lemma 60, one easily sees that for a product of finite type non-amenable groups, $\Gamma=\Gamma_{1}\times\cdots\Gamma_{n}$, the $L^{p}$-cohomology vanishes in every degree $j<n$. It is also not hard to see that in degree $n$ it equals the $L^{p}$-tensor product of the first $L^{p}$-cohomologies of its factors. See [55, p. 252] for a de Rham theoretic proof. This fact motivated Gromov to conjecture a cohomological vanishing below-rank for symmetric spaces and buildings, see [55, p. 253]. Moreover, Gromov conjectured that at the rank the $L^{p}$-cohomology is Hausdorff and that it doesn’t vanish for $p$ large enough, in fact for all $p>1$ in the building case. This conjecture is now fully confirmed.
@@
Over archimedean fields, both the non-vanishing in Statement (2) and the vanishing in Statement (3) are due to Bourdon-Rémy [26]*Theorem A. Over non-archimedean fields, the non-vanishing in Statement (2) is due to [73, Corollary 1.2] and the vanishing in Statement (3) follows by cohomological dimension. In both cases, and both for the groups and their lattices, Statement (1) and the Hausdorffness in Statement (2) is due to [7], being a special case of Corollary 12. For cocompact lattices, the full result follows by Shapiro Lemma and $L^{p}$-induction. ∎
@@
For $\Gamma$ hyperbolic group, following the work of Bourdon-Pajot [25], Nica constructed for every $p$ large enough a non-trivial element in $H^{1}(\Gamma,H^{1}(\Gamma,\ell^{p}\Gamma))$, see [78, Theorem 8]. Using the Lyndon-Hochschild-Serre spectral sequence we have the identification $H^{1}(\Gamma,H^{1}(\Gamma,\ell^{p}\Gamma))=H^{2}(\Gamma\times\Gamma,\ell^{p}\Gamma)$ and we interpret the Nica cocycle accordingly. In particular, we get that $H^{2}(\Gamma\times\Gamma,\ell^{p}\Gamma)\neq 0$.